OpenStudy (anonymous):

The height of an equilateral triangle has length h. What is the length of a side of this triangle? a. h sec [pi]/6. b. hcos pi/6 c. htan pi/6

OpenStudy (anonymous):

 so from the figure, height=h, side=a so, a/h=sec(pi/6) =>a=hsec(pi/6)